A ball is selected at random from a bag containing three red balls, four black balls and five green bars. The first ball is replaced and a second is selected. Find the probability of obtaining two red balls.
1
Expert's answer
2020-03-18T18:23:25-0400

There are a total of "3+4+5=12"

The probability of picking one red ball is "\frac{3}{12}= \frac{1}{4}" . Since the first ball is replaced, the probability of picking a red ball during the second picking is "\frac{1}{4}"

Thus, the probability of picking two red balls with replacement is "\frac{1}{4}×\frac{1}{4}= \frac{1}{16}"